OpenStudy (anonymous):

(5+ (fourth root x)) divided by (fourth root x)

OpenStudy (anonymous):

@bibby you there?

OpenStudy (bibby):

yeah, are we meant to simplify that?

OpenStudy (anonymous):

just rationalize the denominator, then simplify if necessary

OpenStudy (bibby):

not sure how to rationalize the denominator. the best I can do is: \(\huge\frac{5+\sqrt[4]{x}}{\sqrt[4]{x}}\longrightarrow \frac{5}{\sqrt[4]{x}}+\frac{\sqrt[4]{x}}{\sqrt[4]{x}}\longrightarrow\frac{5}{\sqrt[4]{x}}+1\)

OpenStudy (anonymous):

okay so I really just need to know if fourth root x times fourth root x equals x or square root x

OpenStudy (anonymous):

@bibby?

OpenStudy (anonymous):

you want to get the radical out of the denominator you have to multiply the top and bottom by \[\sqrt[4]{x^3}\]